OpenStudy (shubhamsrg):

a chain of uniform mass m and length l is placed on a horizontal table such that 1/4th of its length hangs below the table. Find the work done in pulling whole chain upwards.

OpenStudy (shubhamsrg):

what my attempt was  W = (mg/4)(l/4) = mg/16 but the final ans is 1/2 of this,,what am i missing ?